OpenStudy (abbycross167):

Can someone please help me with a probability math question? please? I will give a medal and fan.

OpenStudy (abbycross167):

Verify that the sum of the experimental probabilities of flipping heads or tails add up to 1. Check both sets of your data. Show your work.

OpenStudy (abbycross167):

Set of data #1 Flip a coin 5 times and record the outcomes here. T= tails H= heads 1. T 2. T 3. H 4. H 5. H

OpenStudy (abbycross167):

Set of data #2 Flip the coin 25 times and keep track of the outcomes. Record them here. T= tails H= heads 1. H 2. H 3. H 4. T 5. H 6. H 7. H 8. T 9. T 10. T 11. T 12. H 13. T 14. H 15. H 16. T 17. H 18. H 19. T 20. T 21. T 22. H 23. H 24. T 25. T

OpenStudy (abbycross167):

Would the answer be like this for the first set of data? 3/5 + 2/5 = 5

jimthompson5910 (jim_thompson5910):

In your data set #1, you have 2 tails. The probability of getting tails, P(T) is P(T) = (# of tails)/(# of flips) = 2/5 similarly, you have 3 heads, so P(H) = (# of heads)/(# of flips) = 3/5 add up the probabilities 2/5 + 3/5 = (2+3)/5 = 5/5 = 1 so that confirms the first data set

jimthompson5910 (jim_thompson5910):

you'll do the same for the second data set

OpenStudy (abbycross167):

oh okay! I get understand now! thank you or the awesome help!

jimthompson5910 (jim_thompson5910):

no problem